3. Find the line y = a + ber that best fits the measurements y=0,1,2,5 at times r=0.1.3, 4.
Answer #1

X 0 1 3 4 Yo 1 2 5

X Y X⋅Y X⋅X
0 0 0 0
1 1 1 1
3 2 6 9
4 5 20 16

Find the sum of every column:

ΣΧ = 8, ΣΥ = 8, ΣΧ. Υ = 27, Σx? = 26

α = ΣΥ.Σ x2 – ΣΧ.Σ ΧΥ ηΣΧ2 - (Σx)? 8. 26 – 8. 27 4. 26 – 82 ο = η ΣΧΥ – ΣΧ. ΣΥ 4. 27 - 8.8 η.ΣΧ2 - (Σ Χ)? 4. 26 - (3) 81.1

y = a + b 2 y = -0.2 + 1.1.